let R be comutetive ring with unity in which each ideal prime then show that R is a field.


1
Expert's answer
2021-06-14T15:50:07-0400

As the zero ideal (0) of R is a proper ideal, it is a prime ideal by assumption.

Hence R=R/{0} is an integral domain.

Let a be an arbitrary nonzero element in R. We prove that a is invertible. Consider the ideal (a2) generated by the element a2. If (a2)=R, then there exists b∈R such that 

1=a2b  as 1∈R=(a2). Hence we have 1=a(ab) and a is invertible. Next, if (a2) is a proper ideal, then

(a2) is a prime ideal by assumption. Since the product "a\cdot a=a^2" is in the prime ideal (a2), it follows that "a\isin (a^2)" . Thus, there exists "b\isin R" such that a=a2b.

Equivalently, we have a(ab-1)=0.

We have observed above that R is an integral domain. As "a\ne 0" , we must have ab-1=0,

and hence ab=1. This implies that a is invertible.

Therefore, every nonzero element of R is invertible.

Hence R is a field.